OpenStudy (anonymous):

find the x-intercepts of the parabola -4.9t^2 + 5t +12.5 Usually we just do this: ___ X ___ = 12.5X-4.9 ---+---= 5 but it doesnt work with decimals...

OpenStudy (anonymous):

Do you know the quadratic formula?

OpenStudy (anonymous):

no but our teacher told us not to use it

OpenStudy (anonymous):

dont you have to find factors of 12.5 X -4, that add to 5? but this doesnt work with decimals

OpenStudy (anonymous):

12.5 X -4.9 *

OpenStudy (anonymous):

Well, if you have learned it, you could try completing the square... but if you're not allowed the quadratic formula you might not have learned this. http://www.purplemath.com/modules/sqrquad.htm Are you saying that you usually factor this? This is factorable (with irrational roots), from what I can tell at first glance.

OpenStudy (anonymous):

Oh, yeah, so: \[ -10 (-4.9t^2 +5t +12.5) = -10(0) \\ 49t^2 -50t -125 = 0 \]

OpenStudy (anonymous):

I multiplied both sides by \(-10\) so it's easier to work with.

OpenStudy (raden):

it can't be factorized

OpenStudy (anonymous):

he just did it ^

OpenStudy (anonymous):

didnt he?

OpenStudy (raden):

just use quadratic formula

OpenStudy (anonymous):

which is?

OpenStudy (raden):

49t^2−50t−125=0 just check the value of discriminant (D) D=b^2-4ac = (-50)^2 - 4(49)(-125) = 2500 + 24.500 = 27,000 sqrt(27,000) isnt a integer numbr

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

cant you factorise it like muntoo

OpenStudy (anonymous):

oi muntoo, wouldnt the second line where you times both sides by ten be this? 49t^2 - 50t - 125 = -10

OpenStudy (anonymous):

how do you get the right side to equall 0?

OpenStudy (raden):

dont forget, not all quadratic equation can be factorized so, to find its solution just use quadratic formula :